# Magnetic phases of the periodic Anderson model in two dimensions

### v1

@@ I Introduction
($\hat{c}_{\bm{k}\sigma}^{\phantom{\dagger}}$) Here, $\hat{c}_{\bm{k}\sigma}^{\dagger}$ ($\hat{c}_{\bm{k}\sigma}$) represents the creation (annihilation) operator for conduction electrons with wave vector $\bm{k}$ and spin $\sigma$. Similarly, $\hat{f}_{j\sigma}^{\dagger}$ ($\hat{f}_{j\sigma}$) denotes the creation (annihilation) operator for localized $f$-electrons at site $\bm{r}_{j}$. The particle number operators for $f$- and $c$-electrons are given by $\hat{n}^{f}_{j\sigma}=\hat{f}_{j\sigma}^{\dagger}\hat{f}_{j\sigma}$ and $\hat{n}^{c}_{j\sigma}$, respectively. The hybridization amplitude between the $f$- and $c$-electron states is denoted by $V$, while $U_{f}$ represents the on-site Hubbard repulsion between $f$-electrons. The average number of $c$- and $f$-electrons per site, denoted as $n_{c}$ and $n_{f}$, respectively, can range from zero to two. The model’s filling is defined as the ratio of the total electron density per site ($n=n_{c}+n_{f}$) to the maximum possible electron occupancy ($n_{\text{max}}=4$). For a comprehensive overview of the fundamental properties of this model, refer to Ref. [6].
@@
In spite of its apparent simplicity, the PAM has no analytic solution even in one dimension unlike the Hubbard model, though exact solutions can be constructed in certain cases [7, 8]. Therefore other techniques have been used to study this model, which often involve approximations. In one dimension the density-matrix renormalization group provides a quasi error-free description of the PAM [9, 10, 11]. At half filling a gapped ground-state emerges with antiferromagnetic correlations between the $f$-electrons mediated by the RKKY interaction. In two dimensions the magnetic properties of the PAM has been extensively explored using quantum Monte Carlo [12], variational Monte Carlo methods [13, 14] as well as dynamical vertex approximation [15]. In three dimensions, the model has been investigated with mean-field theory [16] and, on the Bethe lattice, with dynamical mean-field theory [17, 18]. While dynamical mean-field theory (DMFT) yields exact results in infinite spatial dimensions, it remains uncertain how quantum fluctuations in two and three dimensions modify these findings. The quantum Monte Carlo technique provides reliable results for the half-filled and symmetric ($\varepsilon_{f}=-U_{f}/2$) case, however, moving away from the symmetric case or considering other fillings introduce the sign problem [19], which can significantly complicate or even hinder the simulations. To mitigate the sign problem the constrained path quantum Monte Carlo technique has been successfully used to study the PAM at various fillings [20, 21] but it faces scaling challenges with the system size. Variational techniques have also been widely applied to study the properties of the PAM. These methods either use the Monte Carlo technique [13, 14] or the Gutzwiller approximation [22, 23, 24, 25] to evaluate expectation values. While the former case provides an exact evaluation of the observables and the calculation is only impacted by the variational Ansatz, the latter case involves a further mean-field like approximation as well. Later on, the PAM has been investigated using the density-matrix embedding theory near half filling [26]. Consequently, the analysis of the PAM with techniques that do no rely on rigid Ansätze or uncontrolled approximations is highly warranted.
@@
The goal of this paper is to investigate the ground-state properties of the periodic Anderson model (PAM) on a square lattice at various fillings as a function of the hybridization strength and the $f$-level energy. We employ the infinite projected entangled-pair states (iPEPS) technique [27, 28, 29, 30, 31, 32], which overcomes the aforementioned challenges and has been successfully applied to study frustrated spin systems [33, 34, 35] as well as correlated fermionic systems [36, 37, 38] in two dimensions. Furthermore, we compare our results to those obtained via mean-field theory to emphasize the role of quantum fluctuations. In the symmetric half-filled case, we observe that increasing the hybridization suppresses the antiferromagnetic order, and beyond a critical value, a Kondo singlet phase emerges. Our estimate for the critical point aligns excellently with previous quantum Monte Carlo studies. The antiferromagnetic order persists when the $f$-level energy is increased; however, it is suppressed around $n_{f}\sim 0.5$ occupancy, where a paramagnetic state appears – significantly deviating from the predictions of mean-field theory. Our most striking observation is the emergence of a diagonal stripe state for $n=1.5$ filling hosting a correlated antiferromagnetic order. This state strongly competes with its ferromagnetic counterpart in certain parameter regimes. In general, our studies demonstrate that the iPEPS technique can produce reliable results for challenging parameter regimes and fillings of a correlated two-band model.
@@ II Methods
We use the iPEPS method [27, 28, 29, 30, 31, 32] to study the ground-state properties of the PAM on the square lattice, thus $\varepsilon_{c}({\bm{k}})=-2t(\cos(k_{x})+\cos(k_{y}))$ and we set the hopping amplitude $t$ to $t=1$ and the lattice constant to unity. The iPEPS technique is a variational method that represents the wave function as a network of rank-5 tensors. For a detailed review of the method, see Refs. [27, 28, 29, 30, 31, 32, 39]. The ground state is assumed to be describable using a supercell of tensors that tiles the entire lattice. Each tensor has four auxiliary legs, which connect to neighboring sites, and one physical leg. The accuracy of the approach is controlled by the bond dimension, $D$, of the auxiliary legs. By gradually increasing $D$, the Ansatz can incorporate more entanglement, improving the representation of the ground state. Each tensor contains $dD^{4}$ variational parameters, where $d$ is the dimension of the local Hilbert space. In the periodic Anderson model (PAM), two fermionic degrees of freedom per lattice site are present. To adapt to the algorithm, these are merged into a supersite with a local Hilbert space dimension of $d=16$, allowing the originally developed iPEPS techniques to be directly applied. We optimize the tensors using the fast full update (FFU) algorithm, employing gauge fixing to enhance convergence [40], and compute observables using the corner transfer matrix (CTM) technique [41, 42]. The boundary bond dimension, $\chi$, which determines the accuracy of tensor contractions in the CTM method, was set to $\chi=D^{2}$. Larger $\chi$ values were tested, but their effects were negligible compared to the dependence on $D$ and did not introduce significant quantitative changes to the results. In our simulations, we utilized $\mathrm{U(1)}$ spin and $\mathrm{U(1)}$ charge symmetries to reduce the sizes of the dense tensor blocks. In certain cases, only the $\mathrm{U(1)}$ charge symmetry was applied. The use of both symmetries allowed us to reach bond dimensions up to $D=11$, while simulations using only the $\mathrm{U(1)}$ charge symmetry were limited to $D=9$. Observables were extrapolated to the error-free limit ($D\to\infty$) as a function of $1/D$. For the energies, we employed extrapolation as a function of the normalized cost function with the time step of the imaginary time evolution, $w$, which has been shown to provide higher accuracy. This approach yields smoother energy curves as a function of $w$ compared to $1/D$ [43]. We estimate the extrapolation error as half the distance between the last data point and the extrapolated value.
@@ III.1 Half-filled case
First, we consider the half-filled model at the symmetric point $\varepsilon_{f}=-U/2$, which ensures that the electron density remains for both types of electrons $n_{f}=n_{c}=1$. This case has been investigated in literature with various many-body techniques like quantum Monte Carlo [12] or the dynamical vertex approximation [15] and we choose $U_{f}=4$ so that we can benchmark our results directly against them. Based on the Doniach’s phase diagram, we expect that for small values of the exchange coupling, $J\sim V^{2}/U_{f}$, the RKKY interaction is dominant, which mediates antiferromagnetic fluctuations between the $f$ electrons and leads to magnetic ordering, while for large $J$ the Kondo effect prevails and local singlets are formed. To capture the antiferromagnetic order we use a $2\times 2$ supercell in the iPEPS simulation. The sublattice magnetizations for the conduction ($m_{c}$) and $f$ electrons ($m_{f}$) are defined in the following way:
@@
Figure 1: The sublattice magnetizations, $m_{c,f}$, (upper panel) and local spin correlation between conduction and $f$ electrons (middle panel) as a function of the hybridization in the symmetric ($\varepsilon_{f}=-U_{f}/2$, $U_{f}=4$) half-filled model. For the magnetizations we plotted the mean-field results as well. The bottom panel shows the sublattice magnetization of the $f$ electrons as a function of the inverse bond dimension.
The iPEPS results clearly indicate that the antiferromagnetic phase disappears around $V_{\rm crit}\sim 1.2$, which is in excellent agreement with the quantum Monte Carlo result, $V_{\rm crit}\sim 1.1(1)$ [12], while the dynamical vertex approximation estimates it as $V_{\rm crit}\sim 0.91$ [15]. It is also worth comparing these results to the mean-field predictions. Although the mean-field theory exhibits qualitatively similar behavior, it significantly overestimates the location of the critical point and the magnetization of conduction electrons, indicating that it fails to capture the itinerant magnetization accurately. For small hybridizations, $V\lesssim 0.5$, the conduction electron magnetization $m_{c}$ is indistinguishable from zero, while $m_{f}$ is nearly maximal. The behavior of $m_{c}$ resembles that observed in the Kondo lattice model on the square lattice [45]. Moreover, it can be clearly seen from Fig. 1 (bottom panel) that larger and larger bond dimensions are required as we approach $V_{\rm crit}\sim 1.2$. Above the critical point, the conduction and $f$-electrons form a local singlet. This is evident from the local $\langle\hat{\bm{S}}_{j}^{c}\cdot\hat{\bm{S}}_{j}^{f}\rangle$ correlator, where
@@
and $\vec{\sigma}_{\alpha\beta}$ represents the vector of Pauli matrices acting on the spin indices $\alpha$ and $\beta$. This correlator approaches $-3/4+\langle\hat{n}^{f}_{j\uparrow}\hat{n}^{f}_{j\downarrow}\rangle+\langle\hat{n}^{c}_{j\uparrow}\hat{n}^{c}_{j\downarrow}\rangle$ corresponding to the condition $\langle(\hat{\bm{S}}_{j}^{c}+\hat{\bm{S}}_{j}^{f})^{2}\rangle=0$.
Having established that iPEPS produces reliable results for the symmetric model we now investigate what happens if the $f$-level energy is increased and the system enters the mixed-valence regime.
@@
Fig. 2 clearly shows that the magnetic moments decrease as the $f$ electrons begin to occupy the conduction band. Interestingly, the antiferromagnetic order vanishes around $n_{f}\sim 0.5$ ($\varepsilon_{f}\sim 1.5$), and the system transitions to a paramagnetic state. While the mean-field theory reasonably captures the $f$-level occupancy, it significantly overestimates the stability of the antiferromagnetic state, predicting that the paramagnetic solution becomes energetically favorable only around $n_{f}\sim 0.1$ ($\varepsilon_{f}\sim 3.6$). This poorer performance of the mean-field approach can be attributed to the enhanced charge fluctuations, which are more pronounced in the mixed-valence regime than in the localized moment (symmetric case) regime.
@@
In the remaining part of this paper, we focus on the case where the system is far from half-filling. This scenario has received less attention due to the limited availability of reliable and unbiased methods. Previous studies, including mean-field theory on a cubic lattice [16] and variational calculations based on the Gutzwiller Ansatz on a square lattice [14], predict ferromagnetic ground states over a wide parameter range, particularly when the $f$-level is occupied. Since the ground-state structure is not known a priori, we begin our investigations with a $2\times 2$ unit cell, allowing four independent tensors at each site. By leveraging the $U(1)$ charge and $U(1)$ spin symmetries, we target specific spin sectors of the unit cell corresponding to $n=1.5$ filling. Initially, we examine the parameters $U=4$, $\varepsilon_{f}=-2$, and $V=1$. Our most striking observation is the emergence of a diagonal stripe phase for the $f$ electrons in the iPEPS simulations. In this phase, antiferromagnetic order occurs along the $[1\bar{1}]$ direction, while the sites along the $[11]$ direction exhibit weak magnetization, as shown in Fig. 3(a) (AF state).
@@
This indicates a continuous decrease in magnetization, similar to what occurs in the half-filled case. As the Fermi energy in the conduction band is reduced compared to the half-filled case, the $f$-level becomes empty sooner, cf. the bottom panels of Figs. 2-5. A notable feature common to both the $n=2$ and $n=1.5$ cases is that the magnetizations vanish around $n_{f}\sim 0.5$ in both scenarios. Additionally, we observe that the conduction electrons remain nearly unpolarized, with $m^{c}_{\text{diag}}\sim\mathcal{O}(10^{-3})$.
@@
Having observed that the system at this filling prefers a diagonal stripe pattern, a natural question arises: does this pattern persist in larger unit cells, or can vertical stripes also emerge? To address this, we performed exploratory simulations starting from various random initial states using $2\times 4$ (8 independent tensors) and $4\times 4$ (16 independent tensors) unit cells. In all cases, the diagonal stripe pattern appeared, supporting our initial findings.
@@
Another important issue is the total spin per unit cell, which can vary continuously in an infinite system. To investigate this, we carried out calculations where only $U(1)$ charge conservation was enforced in the iPEPS simulation. This approach, however, limits the maximum bond dimension to $D=9$ due to the larger dense blocks in the tensors. Starting from various random initial states (including states with $\mathcal{O}(1)$ magnetization per unit cell), we found that the total magnetization decreases as the bond dimension increases, eventually approaching zero. This result confirms our assumption that the total magnetization vanishes in the thermodynamic limit. Furthermore, the lowest-energy states consistently exhibited the diagonal stripe pattern.

### v2

@@ I Introduction
Here, $\hat{c}_{\bm{k}\sigma}^{\dagger}$ ($\hat{c}_{\bm{k}\sigma}$) represents the creation (annihilation) operator for conduction electrons with wave vector $\bm{k}$ and spin $\sigma$. Similarly, $\hat{f}_{j\sigma}^{\dagger}$ ($\hat{f}_{j\sigma}$) denotes the creation (annihilation) operator for localized $f$-electrons at site $\bm{r}_{j}$. The particle number operators for $f$- and $c$-electrons are given by $\hat{n}^{f}_{j\sigma}=\hat{f}_{j\sigma}^{\dagger}\hat{f}_{j\sigma}$ and $\hat{n}^{c}_{j\sigma}=\hat{c}_{j\sigma}^{\dagger}\hat{c}_{j\sigma}$, respectively. The hybridization amplitude between the $f$- and $c$-electron states is denoted by $V$, while $U_{f}$ represents the on-site Hubbard repulsion between $f$-electrons. The average number of $c$- and $f$-electrons per site, denoted as $n_{c}$ and $n_{f}$, respectively, can range from zero to two. The model’s filling is defined as the ratio of the total electron density per site ($n=n_{c}+n_{f}$) to the maximum possible electron occupancy ($n_{\text{max}}=4$). For a comprehensive overview of the fundamental properties of this model, refer to Ref. [6].
@@
In spite of its apparent simplicity, the PAM has no analytic solution even in one dimension unlike the Hubbard model, though exact solutions can be constructed in certain cases [7, 8]. Therefore other techniques have been used to study this model, which often involve approximations. In one dimension the density-matrix renormalization group provides a quasi error-free description of the PAM [9, 10, 11]. At half filling a gapped ground-state emerges with antiferromagnetic correlations between the $f$-electrons mediated by the RKKY interaction. In two dimensions the magnetic properties of the PAM has been extensively explored using quantum Monte Carlo [12], variational Monte Carlo methods [13, 14] as well as dynamical vertex approximation [15]. In three dimensions, the model has been investigated with mean-field theory [16] and, on the Bethe and cubic lattices, with dynamical mean-field theory [17, 18, 19, 20]. While dynamical mean-field theory (DMFT) yields exact results in infinite spatial dimensions, it remains uncertain how quantum fluctuations in two dimensions modify these findings. The quantum Monte Carlo technique provides reliable results for the half-filled and symmetric ($\varepsilon_{f}=-U_{f}/2$) case, however, moving away from the symmetric case or considering other fillings introduce the sign problem [21], which can significantly complicate or even hinder the simulations. To mitigate the sign problem the constrained path quantum Monte Carlo technique has been successfully used to study the PAM at various fillings [22, 23] but it faces scaling challenges with the system size. Variational techniques have also been widely applied to study the properties of the PAM. These methods either use the Monte Carlo technique [13, 14] or the Gutzwiller approximation [24, 25, 26, 27] to evaluate expectation values. While the former case provides an exact evaluation of the observables and the calculation is only impacted by the variational Ansatz, the latter case involves a further mean-field like approximation as well. Later on, the PAM has been investigated using the density-matrix embedding theory near half filling [28]. Consequently, the analysis of the PAM with techniques that do no rely on rigid Ansätze or uncontrolled approximations is highly warranted.
@@
The goal of this paper is to investigate the ground-state properties of the periodic Anderson model (PAM) on a square lattice at various fillings as a function of the hybridization strength and the $f$-level energy. We employ the infinite projected entangled-pair states (iPEPS) technique [29, 30, 31, 32, 33, 34], which overcomes the aforementioned challenges and has been successfully applied to study frustrated spin systems [35, 36, 37, 38] as well as correlated fermionic systems [39, 40, 41] in two dimensions. Furthermore, we compare our results to those obtained via mean-field theory to emphasize the role of quantum fluctuations. In the symmetric half-filled case, we observe that increasing the hybridization suppresses the antiferromagnetic order, and beyond a critical value, a Kondo singlet phase emerges. Our estimate for the critical point aligns excellently with previous quantum Monte Carlo studies. The antiferromagnetic order persists when the $f$-level energy is increased; however, it is suppressed around $n_{f}\sim 0.5$ occupancy, where a paramagnetic state appears – significantly deviating from the predictions of mean-field theory. Our most striking observation is the emergence of a diagonal stripe state for $n=1.5$ filling hosting a correlated antiferromagnetic order. This state strongly competes with its ferromagnetic counterpart in certain parameter regimes. In general, our studies demonstrate that the iPEPS technique can produce reliable results for challenging parameter regimes and fillings of a correlated two-band model.
@@ II Methods
We use the iPEPS method [29, 30, 31, 32, 33, 34] to study the ground-state properties of the PAM on the square lattice, thus $\varepsilon_{c}({\bm{k}})=-2t(\cos(k_{x})+\cos(k_{y}))$ and we set the hopping amplitude $t$ to $t=1$ and the lattice constant to unity. The iPEPS technique is a variational method that represents the wave function as a network of rank-5 tensors. For a detailed review of the method, see Refs. [29, 30, 31, 32, 33, 34, 42] and the Appendix. The ground state is assumed to be describable using a supercell of tensors that tiles the entire lattice. Each tensor has four auxiliary legs, which connect to neighboring sites, and one physical leg. The accuracy of the approach is controlled by the bond dimension, $D$, of the auxiliary legs. By gradually increasing $D$, the Ansatz can incorporate more entanglement, improving the representation of the ground state. Each tensor contains $dD^{4}$ variational parameters, where $d$ is the dimension of the local Hilbert space. In the periodic Anderson model (PAM), two fermionic degrees of freedom per lattice site are present. To adapt to the algorithm, these are merged into a supersite with a local Hilbert space dimension of $d=16$, allowing the originally developed iPEPS techniques to be directly applied. We optimize the tensors using the fast full update (FFU) algorithm, employing gauge fixing to enhance convergence [43], and compute observables using the corner transfer matrix (CTM) technique [44, 45]. The boundary bond dimension, $\chi$, which determines the accuracy of tensor contractions in the CTM method, was set to $\chi=D^{2}$. Larger $\chi$ values were tested, but their effects were negligible compared to the dependence on $D$ and did not introduce significant quantitative changes to the results. In our simulations, we utilized $\mathrm{U(1)}$ spin and $\mathrm{U(1)}$ charge symmetries to reduce the sizes of the dense tensor blocks. In certain cases, only the $\mathrm{U(1)}$ charge symmetry was applied. The use of both symmetries allowed us to reach bond dimensions up to $D=11$, while simulations using only the $\mathrm{U(1)}$ charge symmetry were limited to $D=9$. We considered supercells with sizes $2\times 2$, $2\times 4$ and $4\times 4$. Observables were extrapolated to the error-free limit ($D\to\infty$) as a function of $1/D$. For the energies, we employed extrapolation as a function of the normalized cost function with the time step of the imaginary time evolution, $w$, which has been shown to provide higher accuracy (see Appendix for more details). This approach yields smoother energy curves as a function of $w$ compared to $1/D$ [46]. We estimate the extrapolation error as half the distance between the last data point and the extrapolated value.
@@ III.1 Half-filled case
First, we consider the half-filled model at the symmetric point $\varepsilon_{f}=-U_{f}/2$, which ensures that the electron density remains for both types of electrons $n_{f}=n_{c}=1$ (Mott-Hubbard regime). This case has been investigated in literature with various many-body techniques like quantum Monte Carlo [12] or the dynamical vertex approximation [15] and we choose $U_{f}=4$ so that we can compare our results directly to them. Since the quantum Monte Carlo method is free from the sign problem at this point, this serves also as an independent benchmark for the iPEPS technique. Based on the Doniach’s phase diagram, we expect that for small values of the exchange coupling, $J\sim V^{2}/U_{f}$, the RKKY interaction is dominant, which mediates antiferromagnetic fluctuations between the $f$ electrons and leads to magnetic ordering, while for large $J$ the Kondo effect prevails and local singlets are formed. To capture the antiferromagnetic order we use a $2\times 2$ supercell in the iPEPS simulation. The sublattice magnetizations for the conduction ($m_{c}$) and $f$ electrons ($m_{f}$) are defined in the following way:
@@
Figure 1: The sublattice magnetizations, $m_{c,f}$, (upper panel) and local spin correlation between conduction and $f$ electrons (middle panel) as a function of the hybridization in the symmetric ($\varepsilon_{f}=-U_{f}/2$, $U_{f}=4$) half-filled model. For the magnetizations we plotted the mean-field results as well. The bottom panel shows the sublattice magnetization of the $f$ electrons as a function of the inverse bond dimension. The dotted vertical line on the top panel denotes the quantum Monte Carlo estimate for the critical point.
The iPEPS results clearly indicate that the antiferromagnetic phase disappears around $V_{\rm crit}\sim 1.2$, which is in excellent agreement with the quantum Monte Carlo result, $V_{\rm crit}\sim 1.1(1)$ [12], while the dynamical vertex approximation estimates it as $V_{\rm crit}\sim 0.91$ [15]. For small hybridizations, $V\lesssim 0.5$, the conduction electron magnetization $m_{c}$ is indistinguishable from zero, while $m_{f}$ is nearly maximal. For $V=1$ the quantum Monte Carlo study predicts $m_{f}\sim 0.26(5)$, which is again in perfect agreement with the iPEPS value of $m_{f}\sim 0.3$. The behavior of $m_{c}$ resembles that observed in the Kondo lattice model on the square lattice [48]. Moreover, it can be clearly seen from Fig. 1 (bottom panel) that larger and larger bond dimensions are required as we approach $V_{\rm crit}\sim 1.2$, which confirms the increasing correlation length in the system. Above the critical point, the conduction and $f$-electrons form a local singlet. This is evident from the local $\langle\hat{\bm{S}}_{j}^{c}\cdot\hat{\bm{S}}_{j}^{f}\rangle$ correlator, where
@@
and $\vec{\sigma}_{\alpha\beta}$ represents the vector of Pauli matrices acting on the spin indices $\alpha$ and $\beta$. This correlator approaches $-3/4+\langle\hat{n}^{f}_{j\uparrow}\hat{n}^{f}_{j\downarrow}\rangle+\langle\hat{n}^{c}_{j\uparrow}\hat{n}^{c}_{j\downarrow}\rangle$ corresponding to the condition $\langle(\hat{\bm{S}}_{j}^{c}+\hat{\bm{S}}_{j}^{f})^{2}\rangle=0$. Since the sublattice magnetization decreases continuously to zero the transition is of second order between the two phases in agreement with previous studies [12, 48].
It is also worth comparing these results to the mean-field predictions. Although the mean-field theory exhibits qualitatively similar behavior, it significantly overestimates the location of the critical point and the magnetization of conduction electrons, indicating that it fails to capture the itinerant magnetization accurately.
Having established that iPEPS reproduces the quantum Monte Carlo results for the symmetric model we now investigate what happens if the $f$-level energy is increased and the system enters the mixed-valence regime.
@@
Fig. 2 clearly shows that the magnetic moments decrease as the $f$ level empties and the electrons begin to populate the conduction band. Interestingly, the antiferromagnetic order vanishes around $n_{f}\sim 0.5$ ($\varepsilon_{f}\sim 1.5$), and the system transitions smoothly to a paramagnetic state, where the conduction electrons form a singlet state. The overall behavior and the nature of the phase transition agrees well with the previous density-matrix embedding theory investigation [28]. If we compare the iPEPS results to the mean-field theory we can see that the mean-field approach reasonably captures the $f$-level occupancy but it significantly overestimates the stability of the antiferromagnetic state, predicting that the paramagnetic solution becomes energetically favorable only around $n_{f}\sim 0.1$ ($\varepsilon_{f}\sim 3.6$). This poorer performance of the mean-field approach can be attributed to the enhanced charge fluctuations, which are more pronounced in the mixed-valence regime than in the localized moment (symmetric case) regime.
In general, we demonstrated that iPEPS provides accurate results for the half-filled case, therefore we can use it to explore other fillings where the quantum Monte Carlo fails.
@@
In the remaining part of this paper, we focus on the case where the system is far from half-filling. This scenario has received less attention due to the limited availability of reliable and unbiased methods. Since the ground-state structure is not known a priori, we begin our investigations with a $2\times 2$ unit cell, allowing four independent tensors at each site to allow the appearance of more general magnetic structures. By leveraging the $U(1)$ charge and $U(1)$ spin symmetries, we target specific spin sectors of the unit cell corresponding to $n=1.5$ filling. Initially, we examine the parameters $U_{f}=4$, $\varepsilon_{f}=-2$, and $V=1$ (Mott-Hubbard regime) to ensure that the $f$-level is singly occupied. Our most striking observation is the emergence of a diagonal stripe phase for the $f$ electrons in the iPEPS simulations. In this phase, antiferromagnetic order occurs along the $[1\bar{1}]$ direction, while the sites along the $[11]$ direction exhibit weak magnetization, as shown in Fig. 3(a) (AF state). Note that the symmetry-equivalent states can also be found in iPEPS by using either other random initial states or pinning fields in the beginning of the simulation. In what follows, we use the above convention to characterise the state.
@@
This indicates a smooth decrease in magnetization and a phase transition of second order, furthermore the appearance of a paramagnetic state for $\varepsilon_{f}\gtrsim 0$, similar to what occurs in the half-filled case. As the Fermi energy in the conduction band is reduced compared to the half-filled case, the $f$-level becomes empty sooner, cf. the bottom panels of Figs. 2-5. A notable feature common to both the $n=2$ and $n=1.5$ cases is that the magnetizations vanish around $n_{f}\sim 0.5$ in both scenarios. Additionally, we observe that the conduction electrons remain nearly unpolarized, with $m^{c}_{\text{diag}}\sim\mathcal{O}(10^{-3})$, in contrast to the half-filled case.
@@
Having observed that the system at this filling prefers a diagonal stripe pattern, a natural question arises: does this pattern persist in larger unit cells, or can vertical stripes or stripes with other periodicity also emerge? To address this, we performed exploratory simulations starting from various random initial states using $2\times 4$ (8 independent tensors) and $4\times 4$ (16 independent tensors) unit cells. In all cases, the diagonal stripe pattern appeared, supporting our initial findings.
@@
Another important issue is the total spin per unit cell, which can vary continuously in an infinite system. To allow this, we carried out calculations where only $U(1)$ charge conservation was enforced in the iPEPS simulation. This approach, however, limits the maximum bond dimension to $D=9$ due to the larger dense blocks in the tensors. Starting from various random initial states (including states with $\mathcal{O}(1)$ magnetization per unit cell), we found that the total magnetization decreases as the bond dimension increases, eventually approaching zero. This result confirms our assumption that the total magnetization vanishes in the thermodynamic limit. Furthermore, the lowest-energy states consistently exhibited the diagonal stripe pattern.
III.3 Discussion